When I graph y is greater than or equal to x+4 and y is less than 2x-1, there is no overlapping, so I believe I am doing it wrong.
Answer by Menjax(62)

You can use "=" signs:
y=x+4
y=2x-1
2x-1=x+4
x=5
y=9
The point they meet is (5,9)